OpenStudy (kewlgeek555):

Can Someone Tell Me if I Got These Seven Algebra Questions Right? @SolomonZelman @ganeshie8 @mathstudent55 @agent0smith @ranga @DebbieG @nincompoop

OpenStudy (kewlgeek555):

Ryleigh received a $50 gift card to the frozen yogurt shop for her birthday. The shop sells yogurt sundaes for four dollars and yogurt cones for three dollars. How many sundaes and cones can she buy with her card? If x represents the number of yogurt sundaes and y represents the number of yogurt cones, which graph correctly shows the solution to the problem? a. Graph with dashed line, and shaded above boundary line b. Graph with dashed line, and shaded below boundary line c. Graph with solid line, and shaded above boundary line d. Graph with Solid Line, and shaded below boundary line

OpenStudy (kewlgeek555):

I chose B, Graph with dashed line, and shaded blow boundary line. Here is my work: Here is the inequality I created with this word problem:\[4x+3y \le50\]Now here is how I made it into slope-intercept form to figure out the answer:\[4x+3y \le50\]\[4x (- 4x)+3y \le50(- 4x)\]\[3y \le-4x+50\]\[3y(\div 3) \le4x(\div 3)+50(\div 3)\]\[y \le \frac{ -4 }{ 3 }x+\frac{ 50 }{ 3 }\]By this, I can conclude that the graph will have a dashed line and will be shaded below the boundary line. I know because the dashed line is for greater than or equal to and less than or equal to. And shaded below is for less than and less than or equal to.

OpenStudy (kewlgeek555):

Which inequality matches the graph? http://orange.flvs.net/webdav/assessment_images/educator_algebra1_v17/05_05_quiz/05_05_graph_8.gif \[a.~-2x+3y>7\]\[b.~2x-3y<7\]\[c.~-3x+2y \ge7\]\[d.~3x-2y \le7\]Okay, so I know that since this graph has a dashed line, it will not be A or B. I know that it is either C or D. Let me start with C.\[-3x+2y \ge7\]\[-3x(+3x)+2y \ge7(+3x)\]\[2y \ge3x+7\]\[2y(\div2)\ge3x(\div2)+7(\div2)\]\[y \ge \frac{ 3 }{ 2 }x+3.5\]So, based on this equation I know that this is the answer. I know this because this shows a y-intercept of 3.5, a shaded line above the boundary line, and a dashed line, which the graph also shows. So I believe this is the answer.

OpenStudy (kewlgeek555):

Which of the following inequalities match the graph? http://orange.flvs.net/webdav/assessment_images/educator_algebra1_v10/05_05_08.jpg a. -6x + y < 3 b. 6x + y < 3 c. 6x - y < -3 d. The Correct Inequality is Not Listed. I think that the answer is D, The correct inequality is not listed because a dashed line means greater than or equal to and less than or not equal to, and there is none of these type of inequalities listed...

jigglypuff314 (jigglypuff314):

well for the first, I had thought that it would have been a solid line since is "or equal to" so it's including the actual line...

jigglypuff314 (jigglypuff314):

and the other two links have errors...

OpenStudy (kewlgeek555):

Whoa. My teacher taught me this: "Open (>, <) or Closed (_>_, _<_)" turns into "Solid (>, <) or Dashed (_<_, _>_) line.

jigglypuff314 (jigglypuff314):

yeah http://www.sparknotes.com/math/algebra2/inequalities/section1.rhtml " If the inequality is < or > , graph the equation as a dotted line. If the inequality is ≤ or ≥ , graph the equation as a solid line."

OpenStudy (kewlgeek555):

Holy pellet. flutter me. .-.

OpenStudy (kewlgeek555):

So I just. sigh. Okay, can you help me on two that I was confused on, that I didn't answer? They don't requires equation making or solving I believe...

jigglypuff314 (jigglypuff314):

I don't think the links work for the other two... could you try another method?

OpenStudy (kewlgeek555):

The picture links don't work for you? Are you like in Private Mode or something?

jigglypuff314 (jigglypuff314):

it doesn't work even when I'm not in private mode.

OpenStudy (kewlgeek555):

Which of the following inequalities matches the graph? graph of an inequality with a dashed vertical line through the point (2, 0) with shading to the right of the line x < 2 y > 2 y < 2 x > 2

OpenStudy (kewlgeek555):

Which of the following inequalities matches the graph? graph of an inequality with a solid horizontal line through the origin and shading below the line x less than or greater to 0 x greater than or equal to 0 y less than or greater to 0 y greater than or equal to 0

jigglypuff314 (jigglypuff314):

graph of an inequality with a dashed vertical line through the point (2, 0) with shading to the right of the line  so in the horizontal direction (the x direction) there is a dotted line at x = 2 and is shaded to the right (is greater than)